OpenStudy (anonymous):

Find the missing exponent. Set up an equation and solve. : x^4·x^n = x^5 N = missing exponent. Please explain so I can understand.

OpenStudy (freckles):

1 + what =5

OpenStudy (anonymous):

n=1

OpenStudy (freckles):

lol i did it backwards

OpenStudy (anonymous):

when you multiply exponent you add them... x^4*x^n is the same as x^(4+n) so we can say x^(4+n)=x^5 since the bases are the same the exponents need to equal each other... so we can get rid of the bases: 4+n=5 now subtract 4 from both sides... and you get n=1 :)

OpenStudy (anonymous):

n=1 x^4+x^1=x^5

OpenStudy (anonymous):

THANK YOU